A linear regression using risk as the outcome and beds as the predictor produces the following results. Which of the following statements is true based on the below (select all that apply) Linear Fit Risk 3.3735438+0.0070695 Beds Select one or more a. Number of beds is negatively correlated with risk O b. According to the equation, a hospital with 110 beds would have a predicted risk level that is .70695 lower than a hospital with 10 beds c.Number of beds is positively correlated with risk O d. According to the equation, a hospital with 110 beds would have a predicted risk level that is .70695 higher than a hospital with 10 beds
Answer #1

Given equation is ,

Risk = 3.3735438 + 0.0070695*Beds.

Where,   intercept = 3.3735438 , slope = 0.0070695

Slope is positive. So Number of beds is positively correlated with risk.

Predicted risk for hospital with is 110 beds is ,

Risk = 3.3735438 + 0.0070695*110 = 4.151189

Predicted risk for hospital with is 10 beds is ,

Risk = 3.3735438 + 0.0070695*10 = 3.444239

Clearly, a hospital with 110 beds would have a predicted risk level that is 0.70695 (4.151189 - 3.444239 = 0.70695 ) higher than hospital with 10 beds.

So , statements C and D are true based on given linear fit equation.